Sunday, July 15, 2012

Day 49, 7/15/2012 -- Crises, crates, and cables

I realize that I forgot to mention our little liquid helium crisis back on Friday, so I'll mention it here, because the crisis has been summarily averted:

On Friday, we needed to fill the liquid helium tank of the EBEX cryostat. This was no problem, as we had received a full 500-liter dewar about a week and a half prior that would be sufficient to fill EBEX nearly four times over. The only problem is...it wasn't full. In fact, it wasn't anywhere near full. Actually, it turns out it was completely empty. And warm. We found this out Friday morning, and the helium in EBEX would run out on Sunday if we didn't get our fix. And if the helium ran out, the cryostat would start to warm up...and that's not good at all.

So...we let CSBF know, and they set the wheels in motion to see if we could get helium over the weekend. From their liquid helium vendor...in east Texas. They ended up having two CSBF guys drive the 11-12 hours out to the vendor where a full 500L dewar would await them at noon on Saturday, and then drive all the way back so we would have our helium by midnight Saturday. And...so they did! We were able to fill our helium before the situation got really dire -- CSBF really goes out of their way to make sure we're able to get science done.

Aside from helium, today was largely a continuation of yesterday's work. We started adding the cables for the bolo system to the gondola in order that they could be routed in a usable way without risk of them tearing or being sheared by gondola motion. While doing this, I noticed one of the bolo system power cables was noticeably lighter than the others -- it turns out it had been made with the wrong gauge wire (too small), so I set Cynthia to work on making a new cable while Jeff and I worked out the issues routing the other cables. Jeff also mounted the HWP control and readout crate. Kyle spent most of the day debugging commanding issues in Power Crate 1 (with remote help from Ilan over the phone) and seems to have worked out all of the known problems. So with that done, all of the major components of the bolo system were in place on the gondola with the exception of the readout cables.

Elsewhere in the high bay, Andrei (oh, forgot to mention, he was gone and came back on Thursday) and KyleH started setting up for a Bemco test of our liquid cooling system components. In other comings and goings, Franky arrived from McGill late last night to replace Kevin, who will be leaving tomorrow morning. And Kate and daughter left around mid-day today.

I have pictures, but I forgot to grab the memory card from my camera so their posting will have to wait until tomorrow.

UPDATE -- Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120715

No comments:

Post a Comment